# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

### v2

@@ 6 Performance Evaluation of mpfuzz
For the medium setting, with $6$ slots, both baseline B1 and B2 cannot find any DETER exploit in $16$ hours, and baseline B3 can find a DETER exploit in $7.4$ hours. By contrast, mpfuzz finds a DETER exploit in under a minute.
@@
We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$. The tested mempool is configured at $16$ slots. We report the number of exploits found by mpfuzz in a $16$-hour period. Figure 3 presents the result that mpfuzz finds $4096$ exploits in the first $8$ hours on Geth-$v1.10.11$. On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 5.1) within $2$ minutes and does not find anymore in the next $16$ hours. On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour. On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
Figure 2: # exploits found
@@
Figure 3 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours. mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively. On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively. However, the performance of Erigon and Nethermind is much lower than that of Geth. mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively. On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively.
The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool. In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
7 Evaluation of Discovered Attacks
This section shows the evaluation of turning attacks and locking attacks.
7.1 Evaluation of Turning Attacks
7.1.1 Evaluation in Testnet
Setup: We set up our experiment platform by connecting the attacker node to the Goerli testnet. The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
(a) Transaction propagation in six messages
(c) Etherscan screenshot of the blocks generated during the attack on Goerli (txs in six messages)
Figure 4: Mounting $XT_{4}$ attacks on Goerli: Emptying blocks and propagating malicious transactions to the entire network.
To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts. Each account sends $16$ valid pending transactions, followed by a replacement transaction. In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions. The Gas prices of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively. In one setting, we split the $6528$ transactions into six messages (i.e., the first five being $1229$ transactions each and the last one with $384$ transactions). We send these messages in order and wait for $1.5$ seconds between two consecutive messages. For comparison, in another setting, we send the $6144$ valid transactions in one message and the $384$ replacement transactions in another message. Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain. We then turn off the measurement node’s logging and search for the presence of all $6528$ transactions across neighbor nodes in the log. Given a transaction nonce, we report the percentage of neighbor nodes in the testnet that have propagated it to the measurement node.
Table 4: Types of transactions included in the four blocks during the attack in Figure 4c. ‘# private tx’ means the number of private transactions.
# block
# private txs
# public txs
# attack txs
$8040010$
$5$
00
$384$
$8040009$
$78$
$4$
$384$
$8040008$
$1$
00
$384$
$8040007$
$5$
00
$384$
We run experiments of six-message and two-message propagation, repeat each experiment by $3$ times, and report the average result of node percentage in Figure 4a and Figure 4b. Figure 4a shows the transaction propagation with six messages. More than 90% of the testnet nodes have received all of the $15$ valid transactions in a sequence and the replacement transactions. The only exception is the relatively low percentage for the $16$-th transaction of the maximal nonce. By comparison, when sending the transactions in two messages, the propagation fails as shown in Figure 4b where nodes did not propagate any transaction of nonce larger than $5$.
Additionally, Figure 4c shows the generated blocks in the experiment propagating six messages. We took the screenshot from etherscan.io and labeled it in red with information regarding the attack. Before the attack begins, the Goerli testnet normally utilizes $24-47\%$ of the Gas in a block. For ethical consideration, we limit our attack to be short, specifically, four blocks long.
Right after the launch of the attack, in block $8040007$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 4c) is $26.88\%$. The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$. The $6144-384=5760$ child transactions sent in the first round are not included in the block. Other blocks during the attack, namely $8040008$ and $8040010$, are similar. In the third block $8040009$, $73.1\%$ Gas is spent on including $82$ normal transactions as shown in Table 4 not sent by our attacker node. To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node. Furthermore, the private mempool in today’s Ethereum clients [14] has a capability of no limit, which, while being feasible under today’s low private-transaction workloads, can cause resource-exhaustion vulnerability. Notice that we did not discover specific critical nodes to attack as needed by DETER attacks [22].
8 Discussions

### v3

@@ 7.1 Performance
For the medium setting, with $16$ slots, baselines B1, B2, and B4 cannot find any DETER exploit in $16$ hours, and B3 can find Exploit $XT_{3}$ in $7.4$ hours. By contrast, mpfuzz finds the same exploit in under a minute.
The evaluation on different clients is in Appendix C.
7.2 True/False Positive Rates
This section evaluates the true/false positives of the short exploits found by mpfuzz. Recall that mpfuzz runs against a small mempool (MUT) and may introduce false positives, that is, the exploits that satisfy bug oracles in MUT but do not in the actual larger mempool. We formally define true and false positives as follows:
Definition 7.1 (TP/FP exploits)
Consider a short exploit found by mpfuzz on a MUT, $\langle{}st_{0},dc_{0}\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, and a (manually) extended exploit on an actual mempool, $\langle{}st_{0}^{\prime},dc_{0}^{\prime},ops^{\prime}\Rightarrow{}\langle{}st_{n}^{\prime},dc_{n}^{\prime}\rangle{}$.
The pair of exploits constitute a true positive eviction attack w.r.t. $\epsilon$ ($\lambda$), if.f. $st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ st_{0}^{\prime}\cap{}st_{n}^{\prime}=\varnothing\land{}asym_{E}(st_{0}^{\prime},ops^{\prime})<\epsilon$
The pair of exploits constitute a false positive eviction attack w.r.t. $\epsilon$ ($\lambda$), if.f. $st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ (st_{0}^{\prime}\cap{}st_{n}^{\prime}\neq{}\varnothing\lor{}asym_{E}(st_{0}^{\prime},ops^{\prime})>\epsilon)$.
True/false positives on locking attacks can be similarly defined from bug-oracle definitions 4.2 and 4.3.
We evaluate the false positives of the exploits found by mpfuzz under varying $\epsilon$ and $\lambda$. Specifically, we design a two-step experiment: First, given a MUT, we run mpfuzz with varying $\epsilon$ (and $\lambda$). Increasing $\epsilon$ allows mpfuzz to find more short exploits; for each distinct short exploit, we manually extend it to an actual exploit. Second, we evaluate the extended exploits on an actual mempool (in the same local experimental setting as described in § 9).
Table 5: True/false positives (TP/FP) of eviction and locking attacks discovered with varying $\epsilon$ and $\lambda$. ✓means satisfied.
Exploit $\epsilon$ $m=16$ (MUT) Default $m$ TP?
Eq. 1 $asym_{E}$ Eq. 1 $asym_{E}^{\prime}$
Geth-$XT_{1}$ $10^{-4}$ 00 00
Geth-$XT_{3}$09.09.09.09 $0.083$ $0.0002$
Geth-$XT_{6}$125.125.125.125 $0.125$ $0.0003$
Geth-$XT_{2}$ $0.2$ $0.167$ $0.0698$
Geth-$XT_{4}$ $0.23$ $0.208$ $0.0768$
Geth-$XT_{5}$ $0.23$ $0.208$ $0.0768$
Nethermind-$XT_{1}$ $10^{-4}$ 00 00
Nethermind-$XT_{4}$ $0.11$ $0.104$ $0.0008$
Nethermind-$XT_{7}$ $0.36$ $0.355$ $0.0012$
Besu-$XT_{2}$ $0.2$167.167.167.167 $0.0754$
Besu-$XT_{4}$ $0.23$208.208.208.208 $0.0083$
Erigon-$XT_{4}$ $0.23$ $0.208$ $0.0894$
Exploit $\lambda$ $m=16$ (MUT) Default $m$ TP?
Eq. 4 $asym_{D}$ Eq. 4 $asym_{D}^{\prime}$
Reth-$XT_{8}$ $0.34$ $0.34$ $0.015$
OpenEthereum-$XT_{9}$ $0.46$ $0.46$ $0.0439$
Table 5 presents the details of exploits discovered by mpfuzz. For instance, when setting $epsilon$ at $0.0001$ to find eviction attacks on Geth, mpfuzz discovered one short exploit $XT_{1}$ with $asym_{E}=0$ on MUT, which can be extended to an actual exploit on Geth mempool (of $m=6144$) with $asym_{E}^{\prime}=0<0.0001=\epsilon$. This makes it a true positive (marked by ✓in the table). Increasing $\epsilon$ to $0.23$ leads to discovering all six exploits $XT_{1-6}$ on Geth. All these exploits are true positives. The table also includes the results of eviction exploits found on other clients and locking exploits. Overall, the true-positive rate of mpfuzz remains at $100\%$ for $\epsilon\leq{}0.36$ and for $\lambda\leq{}0.46$ across Ethereum clients.
8 Discussions
Root causes of the exploits: We attribute all found exploits $XT_{1-9}$ to four distinct causes: 1) the presence of certain admission patterns (including $XT_{1}/XT_{2}/XT_{3}/XT_{4}/XT_{6}/XT_{7}$). For instance, $XT_{4}$ allows transaction replacement to cause latent overdrafts in place of valid transactions. 2) the absence of certain admission patterns to cause mempool locking (including $XT_{9}$ and $XT_{8}$). For instance, the cause of $XT_{8}$ on Reth is that Reth mempool is a FIFO queue, and it disallows transaction eviction of any kind, which is risky. 3) Inconsistency across multiple admission patterns, notably $XT_{7}$. In $XT_{7}$, the cause is that a mempool allows an eviction and its reversed eviction at the same time. 4) Evadable conditions to context-sensitive admission patterns, including $XT_{3}$ and $XT_{6}$. Both exploits work on Geth, where the mempool triggers the limit of transactions per send under an evadable condition (when the mempool stores more than $py_{3}^{\prime}=5120$ pending transactions).
Mitigation: We propose schemes to mitigate all newly discovered exploits comprehensively. In principle, given an exploit under a known cause, our mitigation design is to negate the cause. For instance, to mitigate $XT_{4}$, we propose declining transaction replacements that cause latent overdrafts. We have implemented this mitigation strategy on Geth, which was merged into the release of Geth $v1.11.4$ [8] . Similarly, to mitigate $XT_{7}$, we propose ensuring the consistency between eviction events and reversed events (e.g., if admitting $tx_{1}$ by evicting $tx_{2}$ is allowed, then admitting $tx_{2}$ by evicting $tx_{1}$ is not allowed). $XT_{1}/XT_{2}$ are mitigated by disallowing eviction of valid transactions by future/latent-overdraft transactions. $XT_{3}/XT_{6}$ are mitigated by removing the triggering condition of transaction sender limit (i.e., making $py_{3}^{\prime}=0$). $XT_{8}$ can be mitigated by re-enabling transaction evictions in certain cases, such as price-based transaction eviction (i.e., a pending transaction arriving later and of higher price can evict another pending transaction of lower price and admitted earlier).
If an exploit relies on multiple causes (e.g., $XT_{3}$ on Causes 1 and 4), negating one cause suffices to mitigate the exploit.
There are more sophisticated cases where one exploit relies on the presence of an admission pattern, and another exploit relies on the absence of the pattern. For instance, $XT_{5}$ relies on the presence of an admission policy named $SAP_{5}$ that turns child transactions into future ones (as in Geth), and the success of $XT_{9}$ relies on the absence of the same policy, that is, $\neg{}SAP_{5}$ (as in OpenEthereum).
To mitigate both $XT_{5}$ and $XT_{9}$, we propose a defense strategy in which the admission decision regarding $SAP_{5}$ is non-deterministic or randomized so that $SAP$ is not always on or off, making it hard for either $XT_{5}$ or $XT_{9}$ to always succeed.
[…]
Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network. For instance, our attack lasts a short period of time, say no more than $4$ blocks produced. We did not test our attack on the Ethereum mainnet. We also mask part of the block numbers in the screenshots (e.g., Figure 3) to prevent detailed attack inspection and reproduction. When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attack by enabling another attack (described above). To prevent introducing new bugs, we did not suggest fixes against turning-based locking ($XT_{5}$, $XT_{6}$) and locking ($XT_{8}$).
9 Conclusion and Future Works
Conclusion: This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles. Running mpfuzz on popular Ethereum clients discovers new mempool-DoS vulnerabilities, which exhibit various sophisticated patterns, including stealthy mempool eviction and mempool locking.
Limitations and future works: The exploit generation in this work is not fully automated. Manual tasks include mempool reduction, mpfuzz setup (configuring $\epsilon$ and symbols), exploit extension, etc. Automating these tasks is the future work.
mpfuzz’s bug oracles neither captures complete dependencies among concrete transactions nor guarantees completeness in finding vulnerabilities. Certified mempool security with completeness is also an open problem.
mpfuzz targets a victim mempool of limited size: the vulnerabilities found in this work are related to transaction eviction, which does not occur in a mempool of infinite capacity. Thus, the mpfuzz workflow cannot find exploits in the mempool of infinite capacity. It is an open problem whether a mempool of large or infinite capacity has DoS vulnerabilities.
10 Acknowledgments
The authors thank USENIX Security reviewers and Kangjie Lu for the discussion in the early stage of this work. Yibo Wang, Wanning Ding, Zhihua Yang, and Yuzhe Tang were partially supported by two Ethereum Foundation academic grants and NSF grants CNS-2139801, CNS-1815814, DGE-2104532. Kai Li is partially supported by NSF grant CNS-2347486 and one Ethereum academic grant.
Appendix A Observations Motivating mpfuzz Design
Tyr’s test oracle detects property violations on post-consensus states. Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low price can be dropped by Ethereum mempools and will not be eventually executed. Besides, they model double-spending transactions as the only type of invalid transaction. In Ethereum, double-spending transactions are of the same nonce. And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
A.1 Further Rationale of Symbolization
We now describe the design rationale of our symbolization technique.
  • The first observation is that real-world mempools admit transactions based on the symbolic value, not concrete value, of nonce and Ether amount. Our design reflects this intuition. For instance, symbols of valid transactions, e.g., $\mathcal{P},\mathcal{C},\mathcal{N}$, are instantiated by mpfuzz to a minimal amount $1$ Ether. As long as the transferred value does not exceed the sender balance, transactions remain valid (i.e., no overdraft). Admission decisions remain the same for two valid transactions despite their difference in nonces or values.
    Likewise, Symbol $\mathcal{F}$ is instantiated by mpfuzz to fixed nonce $n+1$: As long as the nonce is non-consecutive, no matter what specific value it takes (e.g., $3$ or $m+1$), the mempool would deem it as a future transaction and makes the same admission decision.
    As a side note, the nonce of Symbol $\mathcal{F}$ is fixed at a large nonce $n+1$ (we denote it by far future transactions) instead of a small nonce, say $5$ (we denote it by near future transactions), because near future transactions’ nonce could be reconnected as consecutive and lead to duplicated states.
  • The second observation is that restricting eviction/replacement victims to only parent transactions can facilitate finding exploits quickly. Specifically, while any transactions can be evicted/replaced, mpfuzz evicts/replaces only parent transactions; such a strategy can cause maximal damage (i.e., the most child transactions turned) while saving the search space. Our symbol design reflects this idea. For instance, Symbol $\mathcal{R}$ covers any transaction of the same sender and nonce as a transaction in state $st$, but it is instantiated by mpfuzz to only the transaction of nonce $1$ (i.e., replacing a parent transaction).
Appendix B Case Study: How mpfuzz Finds Exploits
We describe how mpfuzz finds an exploit by presenting a case study on finding $XT_{6}$ in the latest Geth $\geq{}v1.11.4$.
Mempool reduction: Recall that a Geth mempool has a capacity of $m^{\prime}=6144$ slots, and its transaction-admission policies are characterized by three essential parameters: admitting up to $py_{1}^{\prime}=1024$ future transactions and limiting up to $py_{2}^{\prime}=16$ pending transactions from any senders when more than $py_{3}^{\prime}=5120$ pending transactions are residing in the mempool. We set up the MUT to run the same codebase or the same admission policy but with smaller parameters: $m=3,py_{1}=1,py_{2}=2,py_{3}=2$ What follows is a description of how mpfuzz find a short exploit on this MUT.
Fuzzing: How mpfuzz automatically finds exploits: Initially, the mempool is filled with $m=3$ normal transactions. That is, the initial symbolized state is $\mathcal{NNN}$. The seed corpus $sdb$ initially contains an empty string. mpfuzz retrieves the empty string and appends to it with different symbols $\mathcal{O}_{0}$, $\mathcal{C}_{0}$, $\mathcal{P}_{0}$. Because of the initial state $\mathcal{NNN}$, only Symbol $\mathcal{P}_{0}$ is feasible. It generates the mutated input $\mathcal{P}$ and instantiates it to a parent transaction of a higher GasPrice than normal transactions (as described in § 5). Sending the input to the mempool gets the transaction admitted, leading to transitioned state $st_{1}=\mathcal{NNP}$. This is a new state that is not in the corpus, and it evicts more normal transactions $\mathcal{N}$ than the previous state $st_{0}$; the input produces positive feedback, and the associated input-state pair $\langle{}P,st_{1}=\mathcal{NNP}\rangle{}$ is added to the corpus.
Figure 6: Snapshot of the mpfuzz state-search tree when finding Exploit $XT_{6}$ on Geth $v1.11.4$.
Next, mpfuzz retrieves from $sdb$ a seed by high energy. According to Table 2, the energy of state $\mathcal{NNN}$ is $\frac{1}{3*3}*0$, and the energy of state $\mathcal{NNP}$ is $\frac{1}{3*2+4}*1=1/10>0$. Hence, seed $\mathcal{NNP}$ is selected. mpfuzz tries input mutation and appends to the selected input $\mathcal{P}$ one of four new symbols, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$ or $\mathcal{P}_{1}$. On state $\mathcal{NNP}$, 1) mutation transaction instantiated from symbol $\mathcal{O}_{1}$ is declined admission. 2) Mutation transaction from $\mathcal{C}_{1}$ is admitted, transitioning the state to $\mathcal{NPC}$, which produces positive feedback and is added to $sdb$. 3) Likewise, transaction $\mathcal{P}_{0}$ is admitted and produces state $\mathcal{NPP}$ of positive feedback; the mutated input is also added to $sdb$. 4) Mutation $\mathcal{P}_{1}$ is admitted but produces an identical state with mutation $\mathcal{P}_{0}$; thus, the mutated input is not added to $sdb$.
Now, there are four seeds in $sdb$: $\mathcal{NNN}(0)$, $\mathcal{NNP}(0)$, $\mathcal{NPC}(1/8)$, and $\mathcal{NPP}(1/12)$. In parentheses are their energy numbers. The seed of the highest energy $\mathcal{NPC}$ is selected ( 1 in Figure 6). mpfuzz then mutates $\mathcal{NPC}$ with $4$ possibilities, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$, and $\mathcal{P}_{1}$, which produce two end states with positive-feedback, that is, $\mathcal{PPC}$ and $\mathcal{PCP}$. They are added to the $sdb$ with energy $\mathcal{PPC}(1/10)$ and $\mathcal{PCP}(1/10)$.
Let’s say $\mathcal{PPC}(1/10)$ is selected ( 2). mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}$ with $7$ mutation transactions, that is, $\mathcal{O}_{1}$, $\mathcal{O}_{2}$, $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, which produces one end state with positive feedback, that is, mutation $\mathcal{C}_{1}$ is admitted and transitions state to $\mathcal{FPC}$. After that, $\mathcal{FPC}$ is the one with the highest energy and is chosen for the next-round fuzzing ( 4). mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}$ with $4$ mutations and produces $2$ state transitions with positive feedback. Upon state $\mathcal{FPC}$, mutation $\mathcal{C}_{1}$ is admitted and leads to state $\mathcal{FEE}$, which satisfies the bug oracle of eviction attacks under $\epsilon=0$. The algorithm then emits the found short exploit: $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ (recall Definition 4.1). The snapshot of the state-search tree is depicted in Figure 6.
Exploit extension: Given the short exploit automatically found on MUT (with $m=3,py_{1}=1,py_{2}=2,py_{3}=2$), the next step is to extend it to a longer exploit functional on the actual Geth mempool (with $m^{\prime}=6144,py_{1}^{\prime}=1024,py_{2}^{\prime}=16,py_{3}^{\prime}=5120$).
Exploit extension requires manual efforts: After identifying exploit $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ is unique, we extend it to the longer exploit by ensuring the same admission event occurs on the actual mempool as on the smaller MUT. In this process, it tries the next transaction of the same sender with the previous one but with an incremented nonce. If it fails, it switches to the next sender. It also tries transaction fees/prices based on measuring the fees of actual normal transactions, which may not be fixed as $3$ in the mpfuzz setting.
Appendix C Additional Eval. of mpfuzz Performance
@@ Appendix B Case Study: How mpfuzz Finds Exploits
We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$. The tested mempool is configured at $16$ slots. We report the number of exploits found by mpfuzz in a $16$-hour period. Figure 8 presents the result that mpfuzz finds $4096$ short exploits in the first $8$ hours on Geth-$v1.10.11$. On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 6.1) within $2$ minutes and does not find anymore in the next $16$ hours. On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour. On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
@@
Figure 8 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours. mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively. On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively. However, the performance of Erigon and Nethermind is much lower than that of Geth. mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively. On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively. The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool. In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
Appendix D Additional Attack Evaluation
D.1 Experiment on a Single Victim Node
Figure 9: Experimental setup
Evaluation settings: Our goal is to evaluate the success rate and cost of different ADAMS attacks on a single victim node. Because some ADAMS attacks are sensitive to the workload of normal transactions, we first collect transaction workloads from the mainnet. Specifically, we instrumented a Geth client (denoted by Geth-m) to log every message it receives from every neighbor. The logged messages contain transactions, transaction hashes (announcements), and blocks. When the client receives the same message from multiple neighbors, it logs it as multiple message-neighbor pairs. We also log the arrival time of a transaction or a block.
Workload collection: We launched a Geth-m node in the mainnet on May 17, 2023, turned on logging for $5$ hours, and collected the transactions propagated to it. We make the collected transactions replayable as follows: We use the account balances and nonces on the mainnet to set up the initial state locally. We then replace the original sender in the collected transactions with the public keys that we generated. By this means, we know the secret keys of transaction senders and are able to send the otherwise same transactions for experiments.
For experiments, we set up three nodes, an attack node sending crafted transactions, a workload node sending normal transactions collected, and a victim node receiving transactions from the other two nodes. The victim node is connected to both the attack and workload nodes. There is no direct connection between the attack node and the workload node. The attacker node runs an instrumented Geth $v1.11.4$ client that can propagate invalid transactions. The victim node runs the tested client. The workload node runs a vanilla Geth $v1.11.4$ client. On each node, we also run a Prysm $v3.3.0$ client at the consensus layer. The experiment platform is denoted in Figure 9. Among the three nodes, only the victim node is staked as a validator and would propose or produce blocks.
In each experiment, we first run the above “attacked” setup. We then run a “regular” setup that excludes the attack node. Under the regular setup, the workload node sends the normal transactions and blocks to the victim node. We compare the experiment results under the attacked setup and regular setup to show the success of the attack.
In the “attacked” setup, we collect the blocks produced and, given a block, we report two metrics: 1) total fees of benign transactions included, 2) total fees of attack transactions included. We also re-run the workload and victim nodes with the “regular” setup. We collect the blocks produced, and, given a block, we report two metrics: 3) total fees of transactions included, and 4) the cost of a baseline spamming attack with 100% success rate. For the latter, given a block, we select the transaction of the highest GasPrice and report the GasPrice multiplied by the Ethereum block Gas limit.
Table 6: Attack success rate and cost (Ether/block)
Clients Exploit Success rate Cost Baseline
Geth $v1.10.25$ $XT_{1}$ $99.80\%$ 00 $11.39$
$XT_{2}$ $99.42\%$ $0.725$ $11.39$
$XT_{3}$ $93.02\%$ $0.0021$ $11.39$
Geth $v1.11.4$ $XT_{4}$ $99.42\%$ $0.806$ $11.39$
$XT_{5}$ $92.65\%$ $0.806$ $11.39$
$XT_{6}$ $94.74\%$ $0.0022$ $11.39$
Erigon $v2.42.0$ $XT_{4}$ $92.53\%$ $1.172$ $17.7$
Nethermind $v1.18.0$ $XT_{4}$ $99.60\%$ $0.0021$ $10.75$
$XT_{7}$ $84.63\%$ $0.20$ $10.75$
Besu $v22.7.4$ $XT_{2}$ $99.63\%$ $1.04$ $17.7$
$XT_{4}$ $99.60\%$ $1.06$ $17.7$